# A framework for fluctuating times and counting observables in stochastic excursions

### v1

@@ VI.3 Cellular sensing
where $\overline{N}$ is the mean number of binding events during the observation, $T_{\text{unbind}}$ is the mean first-passage time from a signaling state to the non-signaling, and $T_{\text{hold}}$ is the mean time before unbinding given a binding event. Immediately, we observe the important role of excursions for the accuracy in cell sensing, since $T_{\text{hold}}=E(\hat{\mathcal{T}})$ is the mean duration of excursions in the sub-set of signaling states, which can be analytically calculated using Eq. (25).
@@
We now consider the uniform ring network studied in Ref. Harvey2023 , where one state is non-signaling and $n-1$ are signaling. All transitions in one direction have rate $r_{+}$ while the others have $r_{-}$. In this model, the energy consumption, measured by entropy production, strongly fluctuates at high-accuracy regimes. Equation (45) allows pinpointing the leading contribution to such large diffusion coefficients. For $n=5$, $r_{+}=4$ and varying values of $r_{-}$, we observe that $\Delta^{2}E(\hat{Q})^{2}/\mu^{3}$ is much larger than the remaining terms when the error is small, as shown in Fig. 6(a). This behavior, which is robust to values of $c$, is caused by the mean entropy production per excursion $E(\hat{Q})$ being orders of magnitude higher than the contributions that include its variance and the covariance with excursion times.
Another insight enabled by applying the excursion framework concerns the interplay between the mean and variance of the excursion duration, whose relation is particularly important in renewal processes, which is the case when there is a single state in subset $A$. In the same setup as in the previous paragraph, we observe in Fig. 6(b) that excursion durations have a sub-Poissonian distribution in high-accuracy regimes where $\text{var}(\hat{\mathcal{T}})<E(\hat{\mathcal{T}})^{2}$, while they become super-Poissonian as estimation errors increase, with $\text{var}(\hat{\mathcal{T}})>E(\hat{\mathcal{T}})^{2}$. This is the case for any value of concentration since the moments of $\hat{\mathcal{T}}$ do not depend on $c$. A super-Poissonian distribution implies more variability in the number of binding events during an observation window, and slower convergences of empirical quantities.
@@ VI.4 Birth-and-death processes
The eigenvectors of $\mathbb{W}_{\!B}=W_{B}-\Gamma$ are the same, and the eigenvalues simply have to be shifted by $-\Gamma$.
@@
Using the decomposition in Eq (90) we find that
$(\psi^{n})_{11}=\sum_{k}\epsilon_{k}^{n}\langle 1|u_{k}\rangle\langle v_{k}|$ (92)
@@
[…]
The integral vanishes when $n$ is odd, which is expected: because of the structure of the process, the excursion must always have the same number of jumps upwards and downwards. When $n$ is even we get
@@
Plugging this back in Eq. (91) yields an analytical formula to evalute the activity distribution:
@@
$P\big{(}\hat{\mathcal{T}}=t\big{)}=2w_{\rm up}w_{\rm dw}e^{-(1+r)w_{\rm dw}t}\frac{I_{1}(2\sqrt{w_{\rm up}w_{\rm dw}}t)}{2\sqrt{w_{\rm up}w_{\rm dw}}t}.$ (104)
@@ VII Conclusion
We have introduced the concept of stochastic excursions in Ref. Fiusa2025 , a framework for characterizing systems far from equilibrium. By focusing on excursions—trajectories from the inactive phase to the active phase and back—we offer a systematic approach to studying the statistical properties of their times and counting observables. This method generalizes existing efforts in sojourn and first passage times, while also enabling the evaluation of meaningful thermodynamic quantities such as dynamical activity, currents, and entropy production, including their fluctuations.
The goal of this article is to consolidate the framework of stochastic excursions and provide a comprehensive characterization of counting observables and related quantities, which are applicable in a wide range of contexts. We derived analytical results for general moments, including moments of the excursion time and counting observables, as well as correlations between excursion duration and counting observables. Explicit formulas were provided for two cases of particular interest: activity (the total number of transitions during an excursion) and counting specific transitions. Furthermore, we demonstrated that, in the long-time limit, the average currents and fluctuations derived from excursion statistics are consistent with steady state results obtained from full counting statistics. This connection bridges the gap between trajectory-level dynamics and ensemble-averaged properties, providing a unified perspective on stochastic excursions. In particular, the decomposition of the diffusion coefficient into components beyond the law of total variance provides a new perspective into the nature of noise in classical stochastic systems. Additionally, we discussed a fluctuation theorem of the exchange type, which holds exactly for individual excursions.
The stochastic excursion framework provides insight into finite-time behavior which is overlooked in steady state analyses. We illustrated this with the example of the absorption refrigerator, where excursions can not only characterize cooling power but also cooling precision by evaluating the probability of failed cooling cycles. In the working parameter range, failed cooling cycles strongly dominate. This characterization is fundamental for practical cooling protocols.
@@ Appendix A Proof of results in the main text
[…]
×$\times P_{\tt exc}(x_{A}\to z_{123\ldots}\to y_{A})\delta\Big{(}t-\sum_{j=1}^{N}\tau_{j}\Big{)}$
[…]
×$\times P_{\tt exc}(x_{A}\to z_{123\ldots}\to y_{A})e^{-iu(\tau_{1}+\ldots+\tau_{N})}$
$=C\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{itu}\sum_{N=1}^{\infty}\langle y_{A}|W_{AB}(\Gamma_{B}+iu)$
×$\times\big{[}W_{B}(\Gamma_{B}+iu)]^{N-1}W_{BA}|x_{A}\rangle$
$=C\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{itu}\langle y_{A}|W_{AB}(iu-\mathbb{W}_{\!B})^{-1}W_{BA}|x_{A}\rangle$
$=C\leavevmode\nobreak\ \langle y_{A}|W_{AB}e^{\mathbb{W}_{\!B}t}W_{BA}|x_{A}\rangle.$
@@
For simplicity we assume a single counting observable, the generalization for any number is straightforward. Start from the stochastic trajectory of an excursion in Eq. (7). The distribution of $\hat{Q}$ and $\hat{\mathcal{T}}$ is
[…]
×$\times P_{\tt exc}\delta\Big{(}\nu_{yz_{N}}+\nu_{z_{N}z_{N-1}}+\ldots+\nu_{z_{2}z_{1}}+\nu_{z_{1}x}-q\Big{)}\delta\Big{(}t-\sum_{j}\tau_{j}\Big{)}.$
@@
$P(q,t)=$ $\ C\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{-i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{-iut}$ (109)
×$\times\sum_{N=1}^{\infty}\sum_{z_{1},\ldots,z_{N}\in B}\int\limits_{0}^{\infty}d\tau_{1}\ldots\int\limits_{0}^{\infty}d\tau_{N}W_{yz_{N}}e^{i\nu_{yz_{N}}\xi}e^{-(\Gamma_{z_{N}}-iu)\tau_{N}}$
×$\times W_{z_{N}z_{N-1}}e^{i\nu_{z_{N}z_{N-1}}\xi}\ldots e^{-(\Gamma_{z_{1}}-iu)\tau_{1}}W_{z_{1}x}e^{i\nu_{z_{1}x}\xi}.$
@@
$P(q,t)=$ $\ C\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{-i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{-iut}$ (110)
×$\times\sum_{N=1}^{\infty}\sum_{z_{1},\ldots,z_{N}\in B}W_{yz_{N}}e^{i\nu_{yz_{N}}\xi}\frac{1}{\Gamma_{z_{N}}-iu}$
×$\times W_{z_{N}z_{N-1}}e^{i\nu_{z_{N}z_{N-1}}\xi}\ldots\frac{1}{\Gamma_{z_{1}}-iu}W_{z_{1}x}e^{i\nu_{z_{1}x}\xi}.$
@@
$P(q,t)=$ $\ C\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{-i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{-iut}$ (111)
×$\times\sum_{N=1}^{\infty}\langle y_{A}|W_{AB\xi}\frac{1}{\Gamma_{B}-iu}\left(W_{\!B\xi}\frac{1}{\Gamma_{B}-iu}\right)^{N-1}W_{BA\xi}|x_{A}\rangle.$
@@
$P(q,t)=$ $\ C\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{-i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{-iut}$ (112)
×$\times\langle y_{A}|W_{AB\xi}\Big{[}iu-\mathbb{W}_{\!B\xi}\Big{]}^{-1}W_{BA\xi}|x_{A}\rangle,$
@@
$P(q,t)=C\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{-i\xi q}\langle y_{A}|W_{AB\xi}e^{\mathbb{W}_{\!B\xi}t}W_{BA\xi}|x_{A}\rangle,$ (113)
which is Eq. (19).

### v2

@@ VI.3 Cellular sensing
[…]
@@
We now consider the uniform ring network studied in Ref. Harvey2023 , where one state is non-signaling and n1n-1 are signaling. All transitions in one direction have rate $r_{+}$ while the others have $r_{-}$. In this model, the energy consumption, measured by entropy production, strongly fluctuates at high-accuracy regimes. Equation (47) allows pinpointing the leading contribution to such large diffusion coefficients. For $n=5$, $r_{+}=4$ and varying values of $r_{-}$, we observe that $\Delta^{2}E(\hat{Q})^{2}/\mu^{3}$ is much larger than the remaining terms when the error is small, as shown in Fig. 7(a). This behavior, which is robust to values of $c$, is caused by the mean entropy production per excursion E(Q^)E($\hat${Q}) being orders of magnitude higher than the contributions that include its variance and the covariance with excursion times.
Another insight enabled by applying the excursion framework concerns the interplay between the mean and variance of the excursion duration, whose relation is particularly important in renewal processes, which is the case when there is a single state in subset $A$. In the same setup as in the previous paragraph, we observe in Fig. 7(b) that excursion durations have a sub-Poissonian distribution in high-accuracy regimes where $\text{var}(\hat{T})<E(\hat{T})^{2}$, while they become super-Poissonian as estimation errors increase, with $\text{var}(\hat{T})>E(\hat{T})^{2}$. This is the case for any value of concentration since the moments of $\hat{T}$ do not depend on $c$. A super-Poissonian distribution implies more variability in the number of binding events during an observation window, and slower convergences of empirical quantities.
@@ VI.4 Birth-and-death processes
Note that here we are considering the steady state regime, and thus the normalization constant $C$ does not depend on initial/final states.
@@
The eigenvectors of $\mathbb{W}_{\!B}=W_{B}-\Gamma$ are the same, and the eigenvalues simply have to be shifted by $\Gamma$.
@@
Using the decomposition in Eq. (88) we find that
${(\psi^{n})_{11}=\sum_{k}\epsilon_{k}^{n}\langle 1|u_{k}\rangle\langle v_{k}|1\rangle}$ (90)
@@
$(\psi^{n})_{11}=(2\sqrt{w_{\rm up}w_{\rm dw}})^{n}\frac{2}{\pi}\int\limits_{0}^{\pi}dk\penalty 10000\ \sin^{2}k\cos^{n}k.$ (92)
The integral vanishes when $n$ is odd, which is expected: because of the structure of the process, the excursion must always have the same number of jumps upward and downward. When $n$ is even we get
@@
Plugging this back in Eq. (89) yields an analytical formula to evaluate the activity distribution:
@@
$P\big(\hat{T}=t\big)=2w_{\rm up}w_{\rm dw}e^{-(1+r)w_{\rm dw}t}\frac{I_{1}(2\sqrt{w_{\rm up}w_{\rm dw}}t)}{2\sqrt{w_{\rm up}w_{\rm dw}}t}.$ (102)
@@ VII Conclusion
We have introduced a framework for characterizing counting observables in stochastic Markov jump processes out of equilibrium in Ref. Fiusa2025 . By focusing on excursions—trajectories from the inactive phase to the active phase and back—we offer a systematic approach to studying the statistical properties of their times and counting observables. This method generalizes existing efforts in sojourn and first passage times, while also enabling the evaluation of meaningful thermodynamic quantities such as dynamical activity, currents, and entropy production, including their fluctuations. Systems whose dynamics have a recurring set of states are naturally phrased in terms of excursions leaving this set and then coming back.
The goal of this article is to provide the tools to readily apply the stochastic framework to a wide range of problems of interest. With that in mind, we consolidate the framework of stochastic excursions and provide a comprehensive characterization of counting observables, their statistics, and related quantities. We derived analytical results for general moments, including moments of the excursion time and counting observables, as well as correlations between excursion duration and counting observables. All results can be written in a relatively simple form that depends only on elements of the stochastic transition matrix. Explicit formulas were provided for two cases of particular interest: activity (the total number of transitions during an excursion) and counting specific transitions. Furthermore, we demonstrated that, in the long-time limit, the average currents and fluctuations derived from excursion statistics are consistent with steady state results obtained from full counting statistics. This connection bridges the gap between trajectory-level dynamics and ensemble-averaged properties. The steady state result can be re-obtained by statistics of counting observables over excursions. In particular, the decomposition of the diffusion coefficient into three distinct components provides a new perspective into the nature of noise in classical stochastic systems. Additionally, we discussed a fluctuation theorem of the exchange type, which holds exactly for individual excursions, showcasing that negative entropy events are exponentially suppressed even at the excursion level. This result was used to derive a thermodynamic uncertainty relation that illustrates the interplay between forward and backward excursion trajectories.
The stochastic excursion framework also provides insight into finite-time behavior which is overlooked in steady state analyses. We illustrated this with the example of the absorption refrigerator, where excursions can not only characterize cooling power but also cooling precision by evaluating the probability of failed cooling cycles. In the working parameter range, failed cooling cycles strongly dominate. This characterization is fundamental for designing practical cooling protocols.
@@ Appendix A Proof of results in the main text
$P\big(\hat{T}=t\big)$ $=\sum_{N=1}^{\infty}\sum_{z_{1},\ldots,z_{N}}\int\limits_{0}^{\infty}d\tau_{1}\ldots\int\limits_{0}^{\infty}d\tau_{N}$ (105)
$\times P_{\tt exc}(x_{A}\to z_{123\ldots}\to y_{A})\delta\Big(t-\sum_{j=1}^{N}\tau_{j}\Big)$
$=\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{itu}\sum_{N=1}^{\infty}\sum_{z_{1},\ldots,z_{N}}\int\limits_{0}^{\infty}d\tau_{1}\ldots\int\limits_{0}^{\infty}d\tau_{N}$
$\times P_{\tt exc}(x_{A}\to z_{123\ldots}\to y_{A})e^{-iu(\tau_{1}+\ldots+\tau_{N})},$
where $P_{\tt exc}$ was defined in Eq. (7). Note that this result is independent of the number of transitions $N$ within one excursion, because we take into account all possible excursion lengths by summing over $N$. Computing the integrals over time provides:
$P(\hat{T}=t)$ $=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{itu}\sum_{N=1}^{\infty}\langle y_{A}|W_{AB}{(\Gamma_{B}+iu)^{-1}}$ (106)
$\times\big[W_{B}{(\Gamma_{B}+iu)^{-1}}]^{N-1}W_{BA}|x_{A}\rangle$
$=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{itu}\langle y_{A}|W_{AB}(iu-\mathbb{W}_{\!B})^{-1}W_{BA}|x_{A}\rangle$
$=C_{x_{A}\to y_{A}}\penalty 10000\ \langle y_{A}|W_{AB}e^{\mathbb{W}_{\!B}t}W_{BA}|x_{A}\rangle,$
which is Eq. (14).
@@
For simplicity we assume a single counting observable, the generalization for any number is straightforward. Assuming that the excursion starts in some state $x_{A}\in A$ and ends in another $y_{A}\in A$, start from the stochastic trajectory of an excursion in Eq. (7). The distribution of $\hat{Q}$ and $\hat{T}$ is
$P(q,t)=$ $\ \sum_{N=1}^{\infty}\sum_{z_{1},\ldots,z_{N}\in B}\int\limits_{0}^{\infty}d\tau_{1}\ldots\int\limits_{0}^{\infty}d\tau_{N}$ (107)
$\times P_{\tt exc}\delta\Big(q-\nu_{y_{A}z_{N}}-\nu_{z_{N}z_{N-1}}-\ldots-\nu_{z_{2}z_{1}}-\nu_{z_{1}x_{A}}\Big)$
$\times\delta\Big(t-\sum_{j}\tau_{j}\Big).$
@@
$P(q,t)=$ $\ C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{iut}$ (108)
$\times\sum_{N=1}^{\infty}\sum_{z_{1},\ldots,z_{N}\in B}\int\limits_{0}^{\infty}d\tau_{1}\ldots\int\limits_{0}^{\infty}d\tau_{N}W_{y_{A}z_{N}}e^{-i\nu_{y_{A}z_{N}}\xi}e^{-(\Gamma_{z_{N}}+iu)\tau_{N}}$
$\times W_{z_{N}z_{N-1}}e^{-i\nu_{z_{N}z_{N-1}}\xi}\ldots e^{-(\Gamma_{z_{1}}+iu)\tau_{1}}W_{z_{1}x_{A}}e^{-i\nu_{z_{1}x_{A}}\xi}.$
@@
$P(q,t)=$ $\ C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{iut}$ (109)
$\times\sum_{N=1}^{\infty}\sum_{z_{1},\ldots,z_{N}\in B}W_{y_{A}z_{N}}e^{-i\nu_{y_{A}z_{N}}\xi}\frac{1}{\Gamma_{z_{N}}+iu}$
$\times W_{z_{N}z_{N-1}}e^{-i\nu_{z_{N}z_{N-1}}\xi}\ldots\frac{1}{\Gamma_{z_{1}}+iu}W_{z_{1}x_{A}}e^{-i\nu_{z_{1}x_{A}}\xi}.$
@@
$P(q,t)=\ C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{iut}$ (110)
$\times\sum_{N=1}^{\infty}\langle y_{A}|W_{AB\xi}(\Gamma_{B}+iu)^{-1}\left[W_{\!B\xi}(\Gamma_{B}+iu)^{-1}\right]^{N-1}W_{BA\xi}|x_{A}\rangle.$
@@
$P(q,t)=$ $\ C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{i\xi q}\int\limits_{-\infty}^{\infty}\frac{du}{2\pi}e^{iut}$ (111)
$\times\langle y_{A}|W_{AB\xi}\left(iu-\mathbb{W}_{\!B\xi}\right)^{-1}W_{BA\xi}|x_{A}\rangle,$
@@
$P(q,t)=C_{x_{A}\to y_{A}}\int\limits_{-\infty}^{\infty}\frac{d\xi}{2\pi}e^{i\xi q}\langle y_{A}|W_{AB\xi}e^{\mathbb{W}_{\!B\xi}t}W_{BA\xi}|x_{A}\rangle,$ (112)
which is Eq. (19). Marginalizing this result over the counting fields immediately provides Eq. (14).
@@
A.4 Excursion thermodynamic uncertainty relation [Eq. (76)]
In this proof, we adapt the steps from Ref. Potts2019 , which derived an uncertainty relation for processes with measurement and feedback. The starting point is the fluctuation theorem of Eq. (75), which we write down as
$P_{x_{A}\to y_{A}}(q,\sigma)=e^{\sigma+\zeta}P_{y_{A}\to x_{A}}(-q,-\sigma),$ (121)
with $\zeta=\log{(C_{x_{A}\to y_{A}}/C_{y_{A}\to x_{A}})}$. Consider the (normalized) auxiliary distribution
$R(q,\sigma)$ $=\dfrac{P_{x_{A}\to y_{A}}(q,\sigma)+P_{y_{A}\to x_{A}}(-q,-\sigma)}{2}$ (122)
$=\dfrac{1+e^{-\sigma-\zeta}}{2}P_{x_{A}\to y_{A}}(q,\sigma),$
whose support is $q,\penalty 10000\ \sigma\in\mathbb{R}$. We now compute a few convenient quantities. We make use of the fluctuation theorem in Eq. (121), the two forms of Eq. (122), the exponential form $\tanh{(x/2)}=(1-e^{-x})/(1+e^{-x})$, and the change of variables $\left\{q,\sigma\right\}\to\left\{-q,-\sigma\right\}$. We start with the first two moments of $\hat{Q}$, where $q$ is the realization of the random variable:
$E(\hat{Q})_{R}=\int dqd\sigma\penalty 10000\ qR(q,\sigma)$ (123)
$=\int dqd\sigma\penalty 10000\ q\dfrac{P_{x_{A}\to y_{A}}(q,\sigma)+P_{y_{A}\to x_{A}}(-q,-\sigma)}{2}$
$=\dfrac{E(\hat{Q})_{x_{A}\to y_{A}}-E(\hat{Q})_{y_{A}\to x_{A}}}{2}.$
And
$E(\hat{Q}^{2})_{R}=\dfrac{E(\hat{Q}^{2})_{x_{A}\to y_{A}}-E(\hat{Q}^{2})_{y_{A}\to x_{A}}}{2}.$ (124)
The variance therefore reads
$\textrm{var}(\hat{Q})_{R}=E(\hat{Q}^{2})_{R}-E(\hat{Q})_{R}^{2}$ (125)
$=\dfrac{\textrm{var}(\hat{Q})_{x_{A}\to y_{A}}+\textrm{var}(\hat{Q})_{y_{A}\to x_{A}}}{2}+\left(\dfrac{E(\hat{Q})_{x_{A}\to y_{A}}+E(\hat{Q})_{y_{A}\to x_{A}}}{2}\right)^{2}.$
Lastly, we shall use the following identity:
$E$ $\left(\left(\hat{Q}-\langle\hat{Q}\rangle_{R}\right)\tanh{\dfrac{\hat{\Sigma}+\zeta}{2}}\right)_{R}=\int dqd\sigma\left(q-\langle\hat{Q}\rangle_{R}\right)$ (126)
$\times\penalty 10000\ \dfrac{1-e^{-\sigma-\zeta}}{2}P_{y_{A}\to x_{A}}(-q,-\sigma)$
$=\dfrac{1}{2}\int dqd\sigma\left(q-\langle\hat{Q}\rangle_{R}\right)P_{x_{A}\to y_{A}}(q,\sigma)$
$-\dfrac{1}{2}\int dqd\sigma\left(q-\langle\hat{Q}\rangle_{R}\right)P_{y_{A}\to x_{A}}(-q,-\sigma)$
$=\dfrac{E(\hat{Q})_{x_{A}\to y_{A}}+E(\hat{Q})_{y_{A}\to x_{A}}}{2}.$
The same calculation applies to the entropy production, where we conclude
$E\left(\left(\hat{\Sigma}+\zeta\right)\tanh{\dfrac{\hat{\Sigma}+\zeta}{2}}\right)_{R}=\dfrac{E(\hat{\Sigma})_{x_{A}\to y_{A}}+E(\hat{\Sigma})_{y_{A}\to x_{A}}}{2}.$ (127)
With all the required identities, we are ready to proceed. First step is to apply the Cauchy-Schwarz inequality:
$\left(\dfrac{E(\hat{Q})_{x_{A}\to y_{A}}+E(\hat{Q})_{y_{A}\to x_{A}}}{2}\right)^{2}$ $=E\left(\left(\hat{Q}-\langle\hat{Q}\rangle_{R}\right)\tanh{\dfrac{\hat{\Sigma}+\zeta}{2}}\right)_{R}$ (128)
$\leq\textrm{var}(\hat{Q})_{R}E\left(\tanh{\left(\dfrac{\hat{\Sigma}+\zeta}{2}\right)^{2}}\right)_{R}.$
The hyperbolic tangent is odd and satisfies $k\tanh{x}\leq\tanh(kx)$ for any $k$ and non-negative $x$. Hence we take $x=(\hat{\Sigma}+\zeta)/2$ and $k=\tanh(\hat{\Sigma}+\zeta)/2$. It then follows
$\tanh{\left(\dfrac{\hat{\Sigma}+\zeta}{2}\right)^{2}}\leq\tanh\left[\tanh\left(\dfrac{\hat{\Sigma}+\zeta}{2}\right)\dfrac{\hat{\Sigma}+\zeta}{2}\right].$ (129)
Note that this result holds irrespective of the sign of $\sigma+\zeta$. Back to Eq. (129), we now get
$\left(\dfrac{E(\hat{Q})_{x_{A}\to y_{A}}+E(\hat{Q})_{y_{A}\to x_{A}}}{2}\right)^{2}$ $\leq\textrm{var}(\hat{Q})_{R}$ (130)
$\times\penalty 10000\ E\left(\tanh\Big[\tanh{\left(\dfrac{\hat{\Sigma}+\zeta}{2}\right)\dfrac{\hat{\Sigma}+\zeta}{2}\Big]}\right)_{R}.$
The argument of the outer tanh is non-negative for any $\{\hat{\Sigma},\zeta\}$, where the function is always concave. Thus we can apply Jensen’s inequality, providing
$\left(\dfrac{E(\hat{Q})_{x_{A}\to y_{A}}+E(\hat{Q})_{y_{A}\to x_{A}}}{2}\right)^{2}$ $\leq\textrm{var}(\hat{Q})_{R}$ (131)
$\times\penalty 10000\ \tanh\left[E\left(\tanh{\left(\dfrac{\hat{\Sigma}+\zeta}{2}\right)\dfrac{\hat{\Sigma}+\zeta}{2}}\right)_{R}\right].$
Finally, recalling Eqs. (123) and (125), and using the exponential form of the hyperbolic tangent, all that is let is elementary algebraic manipulations. We then arrive at
$\dfrac{{\rm var}(\hat{Q})_{x_{A}\to y_{A}}+{\rm var}(\hat{Q})_{y_{A}\to x_{A}}}{\big(E(\hat{Q})_{x_{A}\to y_{A}}+E(\hat{Q})_{y_{A}\to x_{A}}\big)^{2}}\geq\dfrac{1}{e^{[E(\hat{\Sigma})_{x_{A}\to y_{A}}+E(\hat{\Sigma})_{y_{A}\to x_{A}}]/2}-1},$ (132)
which is the thermodynamic uncertainty relation at the excursion level.